OpenStudy (anonymous):

Iris, a biologist needs to find the height of a rare tree. She places a mirror on the ground 42ft from the base of the tree. She moves back, in line with the mirror and the tree, until she can see the reflection of the top of the tree in the mirror.If she is 6ft from the mirroe and her eye is 5ft 4in from the ground, how tall is the tree? help would be greatt<333

OpenStudy (anonymous):

To get rid of the possible confusion of feet and inches I first convert all of the measurements into inches. Ratios: 22.27/60 = x/504 60x = 11224.08 x = 187.068 inches To get the measurement in feet I divide by 12. 187.068 inches / 12 = 15.589 feet. The tree is approximately 15.6 feet tall.

OpenStudy (anonymous):

oh thank you so much!!<33

OpenStudy (anonymous):

Any time!